OpenStudy (anonymous):

Solve the following equation. 6(z-2)+4=10 A.z=2 B.z=3 C.z=-3 D.z=-4 WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (anonymous):

B, 3.

OpenStudy (anonymous):

6 (3-2) = 6 + 4 = 10